# Nullifiers of non-Gaussian cluster states through homodyne measurement

### v1

@@ II Example: Photon subtracted squeezed state
Figure 4: Mean value and standard deviation of the non-Gaussian nullifier estimated from a limited dataset simulated from statistics of an ideal photon-subtracted state. The red line at $0.611$ depicts the Gaussian minimum of the nullifier. Red points correspond to initial squeezing given by $r=0.1$, blue points correspond to $r=0.2$.
@@
The value of the non-Gaussian nullifier of an experimentally prepared cluster state can be estimated from homodyne measurement. When a phase locking is available or a data corresponding to a fixed phase can be post-selected, statistics of four generalized quadratures is sufficient in the discussed example of (16). The photon subtraction in the generation of the cluster state is probabilistic and thus requires stability of the experiment in time. Longer measurement time leads to datasets affected by experimental imperfection. To give an example, as the homodyne detection is phase sensitive, any drifting of the phase affects its precision. Therefore, we analyze the how the size of the dataset affects the precision of the nullifier estimation. We simulate random measurement values, which are drawn from an ideal theoretical probability of measuring $X(\theta)$ in an ideal cluster state. Data are further binned into $1000$ equi-width bins. In each bin, the mean value of simulated data points inside is computed together with their count. The relative frequency of the mean value appearing given by the ratio of counts in a given bin and overall number of data points then approximates the probability. Results from numerical simulation are shown in Fig. 4 for $100-10000$ measurements. The simulation was performed hundred times to determine the spread of results caused by limited sampling from the probabilities. Ten thousand measurement points for each angle are sufficient to certify the non-Gaussianity in the cluster state.
@@ Conclusion
We defined a concept of non-Gaussian nullifiers that enables to certify a successful preparation of non-Gaussian cluster states. We showed that nullifiers defined via Weyl symmetric terms can be easily evaluated from homodyne measurement, an advantageous property as it is the natural measurement of optical cluster states. We present an example of a stabilizer for cluster state created from photon subtracted squeezed states and discuss the effect of imperfect photon subtraction and additional noise in the antisqueezed quadrature. We conclude with a note on requirements on the nullifier precise estimation.
@@ Acknowledgement
VK and PM acknowledge grant 25-17472S of Czech Science Foundation. VK, PM and RF also acknolwedge Horizon Europe Research and Innovation Actions under Grant Agreement no. 101080173 (CLUSTEC). P.M. and R.F. acknowledge project CZ.02.01.010022_0080004649 (QUEENTEC) of EU and the Czech Ministry of Education, Youth and Sport. V.K. and R.F. acknowledges also the Quantera project CLUSSTAR (8C24003) of the Czech Ministry of Education, Youth and Sport and EU. VK acknowledges project IGA- PrF-2025-010 of the Palacký University.
@@
$X(\theta)^{n}=\sum_{k=0}^{n}[\cos(\theta)]^{n-k}[\sin(\theta)]^{k}{n\choose k}:x^{n-k}p^{k}:_{W}.$ (24)
@@
$E=\textrm{diag}(\frac{1}{\cos^{n}(\theta_{1})},\frac{1}{\cos^{n}(\theta_{2})},\frac{1}{\cos^{n}(\theta_{3})}...)$ (27)
@@
$S_{0}=\begin{pmatrix}t\frac{g}{2}+r\frac{1}{2}&0\\ 0&t\frac{1}{2g}+r\frac{1}{2}\\ \end{pmatrix}.$ (30)
@@
$\Pi=1-\outerproduct{0}{0}.$ | . (36)
Finite quantum efficiency of the detector $\eta$ can be simulated by additional losses before an ideal detection Marek et al. (2018). Together with the losses, the corresponding POVM yields [Provaznik thesis]
$\Pi^{\prime}=1-\sum_{k=0}(1-\eta)^{k}\outerproduct{k}{k}.$ | . (37)

### v3

@@ II Example: Photon subtracted squeezed state
Figure 4: Mean value and standard deviation of the non-Gaussian nullifier estimated from a given number of measurements per homodyne phase simulated from statistics of an ideal photon-subtracted state. The red line at $0.611$ depicts the Gaussian minimum of the nullifier. Red points correspond to initial squeezing given by $r=0.1$, blue points correspond to $r=0.2$.
@@
The value of the non-Gaussian nullifier of an experimentally prepared cluster state can be estimated from homodyne measurement. When phase locking is available or if data corresponding to a fixed phase can be post-selected, the statistics of four generalized quadratures is sufficient in the discussed example of (15). The photon subtraction in the generation of the cluster state is probabilistic and thus requires stability of the experiment in time. Longer measurement time leads to datasets affected by experimental imperfections. To give an example, as the homodyne detection is phase sensitive, any phase drift affects its precision. Therefore, we analyze how the size of the dataset affects the precision of the nullifier estimation. We simulate random measurement values, which are drawn from an ideal theoretical probability of measuring $X(\theta)$ in an ideal cluster state. Data are further binned into $1000$ equi-width bins. In each bin, the mean value of simulated data points inside is computed together with their count, creating a histogram. The relative frequency of the mean value appearing given by the ratio of counts in a given bin and the overall number of data points then approximates the probability. Results from numerical simulation are shown in Fig. 4 for 10010000100-10000 measurements. The simulation was performed a hundred times to determine the spread of results caused by limited sampling from the probabilities. Ten thousand measurement points for each angle are sufficient to certify the non-Gaussianity in the cluster state.
III Experiment
In order to illustrate the potential of the nullifier (14) to capture non-Gaussianity in realistic states, we experimentally prepared the single photon state and squeezed single photon (“kitten”) states.
Kitten states at a wavelength of 1550 nm are generated by conditional photon subtraction from a squeezed vacuum state. The squeezed states are produced in an optical parametric oscillator (OPO) containing a 15 mm type-0 periodically poled KTP crystal in a bow-tie ring cavity, pumped at 775 nm by a frequency-doubled laser. The OPO has a bandwidth of 8 MHz, a threshold pump power of $\sim 800$ mW, and an escape efficiency of 97%. For this experiment, the pump power was set to 50 mW.
Photon subtraction is implemented by tapping 3% of the squeezed field on a highly asymmetric beam splitter and detecting a single photon in this arm. To ensure well-defined spectral mode selection, the tapped light passes through a narrow-band Fabry–Perot cavity (78 MHz bandwidth) before coupling into single-mode fiber, where it is further filtered by a 100 GHz DWDM filter, isolating the central degenerate mode at 1550.12 nm. Single-photon detection is performed using a fiber-coupled superconducting nanowire single-photon detector (SNSPD) with  60% detection efficiency and dark count rates of 50–100 Hz. The combined optical filtering and modulation stages yield an overall trigger channel efficiency of  20%. All optical cavities and measurement phases are actively stabilized using feedback control. During data acquisition, bright locking beams are momentarily blocked to prevent degradation of squeezing and saturation of the photon detector, employing a lock–measure cycle at a 100 Hz repetition rate.
The single photon state is prepared in a very similar way, with two main differences: The crystal is changed to one with type-2 phase matching, in order to generate polarisation non-degenerate photon pairs, and the output is split on a polarising beam-splitter instead of a 97/3 splitter. With this crystal, the pump threshold is rather high – around 20 W – so even with the pump power of 100 mW, the multi-pair production rate is negligible.
The conditionally prepared states are characterized via time-resolved balanced homodyne detection. The homodyne detector combines a balanced beam splitter with a balanced detector (40 MHz bandwidth, 15 dB clearance) and achieves an overall detection efficiency of  88%. A strong local oscillator allows measurement of quadratures at controlled phases, with 10,000 traces acquired (5,000 for the single photon) at each 30 phase interval between 0 and 150. Each trace is recorded in a 1 µs window centered on the trigger click, sampled at 500 MS/s to resolve the OPO’s temporal correlations. Quadrature values are extracted by integrating each trace against a mode function matched to the OPO’s temporal response and the detection bandwidth. This choice maximizes the reconstructed Wigner function negativity.
From the phase-resolved quadrature distributions, the density matrix and Wigner function of the generated state are reconstructed using a Maximum Likelihood algorithm. We do not correct for any losses or inefficiencies. This procedure yields the phase-space representation of the heralded non-Gaussian states, as illustrated in figure 5. Both states have Wigner function minima of W(0,0)=0.11W(0,0)=-0.11.
Figure 5: Wigner functions of the experimentally prepared single photon state and squeezed single photon state.
The expectation value of the nullifier was obtained with eq. (16) as $0.3976(36)$ for $g=0.9752$ for the single photon state and $0.4810(20)$ for the squeezed single photon state with $g=1.6978$. Both states have the expectation value below the Gaussian threshold of $0.611$. The derivation of the uncertainty is described in more detail in Appendix C.
@@ Conclusion
We introduced the concept of non-Gaussian nullifiers as a practical tool for certifying the successful preparation of non-Gaussian cluster states. We showed that nullifiers constructed from Weyl symmetric operators can be directly evaluated from homodyne measurement, which is particularly advantageous as homodyne detection is the natural detection method for optical cluster states. As a concrete example, we considered a nullifier for a cluster state generated from photon-subtracted squeezed states and analyzed the impact of experimental imperfections, including imperfect photon subtraction, excess noise in the anti-squeezed quadrature, and finite precision in nullifier estimation. More generally, this framework extends Gaussian nullifier techniques into the non-Gaussian regime and provides a practical route toward verifying resource states for universal continuous-variable quantum computation. It is applicable to a wide class of non-Gaussian states, including cubic phase and GKP states, and offers a scalable alternative to full state tomography for future photonic quantum processors. Finally, we demonstrated experimentally how the proposed non-Gaussian nullifier can be used to identify photon-subtracted squeezed states.
@@ Acknowledgement
VK and PM acknowledge grant 25-17472S of Czech Science Foundation. JSNN, PM, RF, ULA and VK also acknolwedge Horizon Europe Research and Innovation Actions under Grant Agreement no. 101080173 (CLUSTEC). P.M. and R.F. acknowledge project CZ.02.01.010022_0080004649 (QUEENTEC) of EU and the Czech Ministry of Education, Youth and Sport. JSNN, ULA, V.K. and R.F. acknowledges also the Quantera project ClusSTAR (8C24003) of the Czech Ministry of Education, Youth and Sport, the Innovation Fund Denmark and EU. JSNN and ULA acknowledge grant no.1063-00046A from Innovation Fund Denmark (PhotoQ). JSNN and ULA acknowledge project bigQ no. DNRF0142 from Danish National Research Foundation. JSNN and ULA acknowledge EU ERC project (ClusterQ no. 101055224). VK acknowledges project IGA- PrF-2025-010 of the Palacký University. VK acknowledges support from the F.R.S.–FNRS under project CHEQS within the Excellence of Science (EOS) program.
@@
$X(\theta)^{n}=\sum_{l=0}^{n}[\cos(\theta)]^{n-l}[\sin(\theta)]^{l}{n\choose l}:x^{n-l}p^{l}:_{W}.$ (23)
@@
[…]
@@
$S_{0}=\begin{pmatrix}t^{2}\frac{g}{2}+r^{2}\frac{1}{2}&0\\ 0&t^{2}\frac{1}{2g}+r^{2}\frac{1}{2}\\ \end{pmatrix}$. (29)
@@
$|.\Pi=1-\outerproduct{0}{0}$. (35)
Finite quantum efficiency of the detector $\eta$ can be simulated by additional losses before an ideal detection. Together with the losses, the corresponding POVM yields Marek18
$|.\Pi^{\prime}=1-\sum_{k=0}(1-\eta)^{k}\outerproduct{k}{k}$. (36)
IV Appendix C: Precision estimation of the nullifier mean value derived from homodyne data
The mean value of the nullified can be computed as
$\expectationvalue{O}=\sum_{k}c_{k}\expectationvalue{X(\theta_{k})^{n_{k}}}.$ (37)
Having a homodyne measurement sample $\{X_{i}(\theta_{k})\}$ of size $N$, the moments can be estimated as
$\expectationvalue{X(\theta_{k})^{n_{k}}}=\frac{\sum_{i}X_{i}(\theta_{k})^{n_{k}}}{N},$ (38)
the variance of the estimator can be found as
$\sigma^{2}_{k}=\frac{\sum_{i}X_{i}(\theta_{k})^{2n_{k}}-\left(\expectationvalue{X(\theta_{k})^{n_{k}}}\right)^{2}}{N-1}.$ (39)
In a similar manner covariances between potentially dependent moments is
$\begin{split}&c_{kj}=\\ &\frac{(\sum_{i}X_{i}(\theta_{k})^{n_{k}}-(\expectationvalue{X(\theta_{k})^{n_{k}}}))(\sum_{i}X_{i}(\theta_{j})^{n_{j}}-(\expectationvalue{X(\theta_{j})^{n_{j}}}))}{N-1}.\end{split}$ (40)
The overall error can then be computed by
$\sigma^{2}=\sum_{k}c_{k}^{2}\sigma^{2}_{k}+\sum_{kj}c_{kj}$ (41)